Circular No. 8173

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003gu IN UGC 12331 Further to IAUC 8103, M. Armstrong reports the discovery of an ap-

parent supernova (mag 17.2) on unfiltered CCD images taken on Aug. 3.051 and 3.965 UT. The new object is located at α = 23^{h}02^{m}59^{s}.45, δ = +34^{o}43′19′′.6, which is ≈ 1′′ east and 18′′.5 south of the center of UGC 12331. Nothing was visible at this position on Armstrong’s image taken on July 21 (limiting mag 18.5) and on Palomar Sky Survey red or blue plates.

SUPERNOVAE 2003gh, 2003gi, 2003gr M. Hamuy and N. Morrell, Carnegie Observatories; J. Arias, Observa-

torio de La Plata; and R. Barba, Universidad de La Serena, report that a spectrum (range 390–900 nm) of SN 2003gr (cf. IAUC 8169), obtained on July 31.31 UT with the Las Campanas Observatory Baade 6.5-m telescope (+ Boller & Chivens spectrograph), shows it to be a type-Ia supernova. The spectrum bears strong resemblance to that of SN 1994D taken 11 days after maximum. The minimum of the Si II 635.5-nm line yields an expansion velocity of −10200 km/s (adopting the NED host-galaxy recession velocity of 7691 km/s). A spectrum of SN 2003gh (cf. IAUC 8159), taken on July 31.44, also shows it to be a type-Ia supernova. The spectrum is similar to that of SN 1994D taken 38 days past maximum. A spectrum of SN 2003gi (cf. IAUC 8161), taken on July 31.32, shows it to be a type-Ia supernova. The spectrum is similar to that of SN 1994D taken 76 days past maximum.

SUPERNOVA 2003gs IN NGC 936 Hamuy also reports, with L. Gonzalez (Universidad de Chile), the fol-

lowing preliminary CCD magnitudes for SN 2003gs (cf. IAUC 8171) from data obtained with the Las Campanas Observatory Swope 1-m telescope: July 30.33 UT, B = 14.12, V = 13.46, I = 13.31; July 31.27, 14.23, 13.48, 13.30. These observations suggest that SN 2003gs is a few days past V maximum. The B − V colors (0.66 and 0.73) at these two epochs is much redder than the color displayed by normal type-Ia events near maximum light (B − V = 0) but is similar to that of the subluminous SN 1991bg (Leibundgut et al. 1993, A.J. 105, 301). A spectrum (range 390–900 nm), obtained as above by N. Morrell on July 31.33 with the Baade telescope, confirms that SN 2003gs is spectroscopically peculiar (cf. IAUC 8171, 8172). The lack of interstellar Na I D lines suggests that SN 2003gs is not significantly reddened by dust.

2003 August 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green